OpenStudy (anonymous):

what is the inverse of f(x)=3x-5 PLEASSEEEE HELLPPPP its from the project on flvs

OpenStudy (anonymous):

im definitely new on here lol

OpenStudy (anonymous):

Inverse is just doing what happened to the function but in reverse order and opposite. First you multiplied by three then subtracted five. so the invers is (x+5)/3

OpenStudy (anonymous):

so i would use that function and solve it the opposite way?

OpenStudy (anonymous):

A common way is solving for x then switching out the x and y. But I just think of it like the way before

OpenStudy (anonymous):

would it be x+2=y?

OpenStudy (anonymous):

im probably wrong lol

OpenStudy (anonymous):

No, y=3x-5. Add five then divide by three and you get (y+5)/3=x. Then you switch your x and y and get (x+5)/3. That's the other to think about it

OpenStudy (anonymous):

OHHHHHH thank you:) helped alot

OpenStudy (anonymous):

One last question @Prominence, becaus eit is opposite would the function end up being (y+5)/3=x or (x+5)/3=y

OpenStudy (anonymous):

since it has to be the opposite of y=3x-5